undeadknight26 (undeadknight26):

Derive the equation of the parabola with a focus at (4, −7) and a directrix of y = −15. Put the equation in standard form.

undeadknight26 (undeadknight26):

@aaronq @Alchemista @SithsAndGiggles @sourwing @dan815 @freckles @ganeshie8 @goatdude101 @hhelpplzzzz @KamiBug @LazyBoy @zepdrix @cwrw238 @mathstudent55 @Mimi_x3 @mathmath333 Please help ;-;

OpenStudy (anonymous):

Hmm, so the distance between the point and the line are always equal

OpenStudy (anonymous):

I mean the point and parabola and the directrix and parapola.

undeadknight26 (undeadknight26):

f(x) = (1/12)x^2 - 8x + 11 f(x) = (1/16)x^2 - 8x - 10 f(x) = (1/16)x2 - (1/2)x + 11 f(x) = (1/16)x^2 - (1/2)x - 10

OpenStudy (anonymous):

wio can you check my answer after this?

OpenStudy (anonymous):

\[ \sqrt{(y-(-15))^2} = \sqrt{(x-4)^2+(y-(-7))^2} \]Square both sides: \[ (y+15)^2=(x-4)^2+(y+7)^2 \]Simplify:\[ y^2+30y+225=x^2-8x+16 +y^2+14y+49\\ 16y = x^2-8x-160\\ y = \frac{x^2-8x-160}{16} \]

OpenStudy (anonymous):

I used a long approach, there is another approach involving \[ 4py= x^2 \]Though I don't completely remember it.

undeadknight26 (undeadknight26):

So its D?

OpenStudy (anonymous):

Well, let \(x=4\). See what the value of \(y\) is. If the distances check out, you have your solution

OpenStudy (anonymous):

please check wio

undeadknight26 (undeadknight26):

I got -26...